OpenStudy (sierra_tibbetts):

Simplify square root of 2 over cube root of 2. 2 to the power of 1 over 6 2 to the power of 1 over 3 2 to the power of 3 over 2 2 to the power of 5 over 6

OpenStudy (sierra_tibbetts):

\[\frac{ \sqrt{2} }{ \sqrt[3]{2} }\]

OpenStudy (sierra_tibbetts):

@ShadowLegendX